Date: Tue, 8 Jun 2004 04:46:19 -0700 (MST)
From: "Ettore Pedretti, F1VGM" <>
Subject: end of run report

The last two nights were characterised by bad seeing and nearly useless 
data because of that. In addition John discovered oscillations of SD2 in 
some of the data. These oscillations preclude calibration. The 5 mv 
oscillation seen on the oscilloscope in random points of the track is not 
acceptable and indicates problems on the recorded data. Also when I tried 
to reduce pressure the error signal went up and we lost the position of 
SD2. This happens for a pressure slightly low than 55. All is well 
documented in the observing log.

Other big problem during the run was the pointing model of telescope C. We 
hardly found a star in the field of the acquisition camera. This is a 
really time-wasting problem.

Other problem found is that the piezo search in the star-tracker does not 
seem to work when acquiring faint stars. The star-tracker does can lock the 
stars but the piezo search just does not seem to detect the star. Is there 
a threshold to set somewhere?

CPU2 is consistently not booting during a VME cold start. We have to 
reboot in order to be able to download the software. Is it an indication 
of worse times to come? Do we need a spare?

That's all for now.

Rebecca Ming and ep